OpenStudy (anonymous):

divide and rationalize 3+sqrt(18)/ 1+sqrt(8)

OpenStudy (anonymous):

hey, how's it going?

OpenStudy (anonymous):

hi can you help me?

OpenStudy (anonymous):

sure, first do you know how to simplify sqrt(18) and sqrt(8)?

OpenStudy (anonymous):

yes i simplified the problem to \[3+3\sqrt{2}/1+2\sqrt{2}\]

OpenStudy (anonymous):

good!

OpenStudy (anonymous):

when you have a square root in the denominator, to clean it up you have to multiply by something called the conjugate, which basically just means you change the sign of the square root term

OpenStudy (anonymous):

i did that too.

OpenStudy (anonymous):

so for this case, you would multiply it by (1-2sqrt(2))/(1-2(sqrt(2))

OpenStudy (anonymous):

\[1-2\sqrt{2}\] right?

OpenStudy (anonymous):

right, so did you multiply the top and bottom by that?

OpenStudy (anonymous):

i did but i erased it.

OpenStudy (anonymous):

but just foil right? i get the top part but the bottom confuses me.

OpenStudy (anonymous):

what confuses you about the bottom part? what did you get for the top?

OpenStudy (anonymous):

i got \[3-6\sqrt{2}+6\sqrt{2}-12\] simplified to -9

OpenStudy (anonymous):

WHat hold on somyjing looks rwong

OpenStudy (anonymous):

thats what i figured i got everything till the foiling part

OpenStudy (anonymous):

bob how did you do this

OpenStudy (anonymous):

2sqrt(2)*3sqrt(2) is what?

OpenStudy (anonymous):

\[6\sqrt{2}\]

OpenStudy (anonymous):

well no, that is 12, but you got that part right, your mistake was 3sqrt(2) * 1 i believe

OpenStudy (anonymous):

oh your right... i thought i was going to be able to eliminate the roots in the middle...

OpenStudy (anonymous):

you want to eliminate the roots on the bottom, that's why are we doing this. you don't have to eliminate the roots on the top.

OpenStudy (anonymous):

\[3-6\sqrt{2}+3\sqrt{2}\] -12 right?

OpenStudy (anonymous):

looks good for the top

OpenStudy (anonymous):

i kept getting 1-8

OpenStudy (anonymous):

looks good, so what's your end answer?

OpenStudy (anonymous):

hold on i haven't solved it yet, thought you were going to say i was wrong.

OpenStudy (anonymous):

where's your confidence? :)

OpenStudy (anonymous):

not in my math capabilities thats for sure haha. \[-9-3\sqrt{2}/-7\]

OpenStudy (anonymous):

okay so the numbers are right but my text book says all the nubmer are positive.

OpenStudy (anonymous):

looks good, don't forget parentheses if you write it like you typed it, and you can multiply by -1/-1 and get rid of all the negatives if you want

OpenStudy (anonymous):

parentheses? the ones used to foil? and okay thank you soo much your very helpful.

OpenStudy (anonymous):

i was just saying the way you typed your answer, could be -9 - (3sqrt(2)/-7) which is not the same thing